OpenStudy (anonymous):

Cashews cost $5 per pound and peanuts cost $1.50 per pound. How many pounds of cashews should be mixed with peanuts to get a mixture of 32 pounds that will cost $4 per pound? Round to the nearest tenth

OpenStudy (anonymous):

Total Cost = Amount of pounds x Cost per Pound = 32 x 4 = 128 Let Pounds of Cashew = C. Total Pounds = Pounds of Cashew + Pounds of Peanuts 32 = C + Pounds of Peanuts Pounds of Peanuts = 32 - C Total Cost = 128 = Cost of Peanuts + Cost of Cashews 128 = (32 - C)(1.5) + (C)(5) 128 = 48 - 1.5C + 5C 128 = 48 +3.5C C = ?

OpenStudy (anonymous):

18.4 10.3 22.9 9.1 these are the answers